# Generative AI models enable efficient and physically consistent sea-ice simulations

### v1

@@ Appendix A RMSE Score cards
Appendix B Mathematical framework for surrogate modelling
@@
Our goal is to develop a powerful surrogate model that forecasts sea-ice states with a 12-hour lead time, $\mathbf{x}_{t+12\,\text{h}}$, given some initial states $\mathbf{x}_{t}$ and external forcings $\mathbf{F}_{t:t+12\,\text{h}}$ provided at initial time $t$ and 12 hours later, $t+12\,\text{h}$. As a reminder, the predictive variables are the sea-ice thickness, the concentration, the damage, the two velocity components, and the snow-on-ice thickness. Our model predicts the tendency within the 12 hours,
$\mathbf{x}_{t+12\,\text{h}}\approx\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}}=\mathbf{x}_{t}+\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\mathbf{F}_{t:t+12\,\text{h}}),$ (15)
@@
with $\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}}$ as the model prediction. The tendency model $\mathcal{M}_{\boldsymbol{\theta}}(\cdot)$ is based on a deep neural network with its learnable weights and biases $\boldsymbol{\theta}$. Here, we train two different types of models: a deterministic model [12] and a generative model based on flow matching [23, 25, 24].
@@
To train the models, we sample data triplets from a training dataset $(\mathbf{x}_{t},\mathbf{x}_{t+12\,\text{h}},\mathbf{F}_{t:t+12\,\text{h}})\sim\mathcal{D}$ and minimise the loss functions using variants of stochastic gradient descent. Once trained, we employ the models as outlined in Eq. (15) to propagate the initial states forward and to generate a 12-hour lead time prediction. For predictions extending beyond 12 hours, we utilise the model autoregressively, wherein the preceding prediction serves as the input for the subsequent time step, with the forcings being updated accordingly.
@@
$\mathbf{z}=\frac{\mathbf{x}_{t+12\,\text{h}}-\mathbf{x}_{t}}{\boldsymbol{\sigma}_{\textrm{clim}}}.$ (16)
@@ B.1 Deterministic model
$\mathbf{x}_{t+12\,\text{h}}\sim\mathcal{N}\left(\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}},\boldsymbol{\sigma}_{\textrm{clim}}^{2}\mathbf{I}\right),$ (17)
@@
$\mathcal{L}_{\textrm{det}}(\boldsymbol{\theta})=\mathbb{E}_{\mathcal{D}}\left[\mathbf{w}\lVert\mathbf{x}_{t+12\,\text{h}}-\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}}\rVert^{2}_{2}\right],$ (18)
with $\mathbf{w}=1/{\boldsymbol{\sigma}_{\textrm{clim}}^{2}}$ as weighting vector. This loss function is commonly used for other surrogate models [72, 73, 12] and serves as a baseline for our generative model.
@@
$\mathcal{L}_{\textrm{det}}(\boldsymbol{\theta})=\mathbb{E}_{\mathcal{D}}\left[\lVert\mathbf{z}-\widehat{\mathbf{z}}_{\boldsymbol{\theta}}\rVert^{2}_{2}\right].$ (19)
@@ B.2.1 Probabilistic predictions with flow matching
$\mathbf{z}_{0}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}).$ (20)
@@
$\frac{\mathrm{d}\mathbf{z}_{\tau}}{\mathrm{d}\tau}=\mathbf{u}(\mathbf{z}_{\tau}),$ (21)
@@
$\mathbf{z}_{\tau}=\tau\mathbf{z}_{1}+(1-\tau)\mathbf{z}_{0},$ (22)
@@
$\mathbf{u}=\mathbf{z}_{1}-\mathbf{z}_{0}.$ (23)
Since $\mathbf{z}_{1}$ is targetted and the analytical velocity therefore unknown during prediction, we train the neural network with its parameters $\boldsymbol{\theta}$ to approximate the true velocity, $\mathbf{u}\approx\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})$. By conditioning the approximation on the initial states and forcings, the neural network learns to sample from the conditional distribution [74].
@@
$\mathcal{L}_{\textrm{FM}}(\boldsymbol{\theta})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\lVert\mathbf{u}-\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z}_{\tau},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})\rVert^{2}_{2}\right].$ (24)
Once trained, we can employ the neural network as velocity approximation for the ODE, Equation (21). By initialising with Equation (20) and integrating the ODE in pseudo time from $\tau=0$ to $\tau=1$, we produce a sample from the target distribution, conditioned on the initial states and forcings. By drawing different initialisations, an ensemble of predictions can be produced from the same conditional information. In practise, we perform the integration with a second-order Heun scheme [63] in 20 steps, where the discrete steps are scheduled with a shifted sigmoid function, as shown in Fig. 9b.
Notably, if projected back into physical space, the generative flow is initialised by a sample from the prior distribution around the initial conditions $\mathcal{N}(\mathbf{x}_{t},\boldsymbol{\sigma}^{2}_{\text{clim}}\mathbf{I})$ and maps to a sample from the predictive distribution for $\mathbf{x}_{t+12\,\text{h}}$. Although a deterministic forecast, trained with Equation (18), could also be used to define the prior [29, 48], previous work [35] has demonstrated negligible performance difference, and we therefore employ a persistence forecast as prior for simplicity. Since the integration of the ODE maps from the initial conditions to the prediction, it works similarly to classical geophysical models [27], integrating in pseudo time instead of real time. Hence, we can make use of established numerical tools to render generative flow models more efficient and to scale them up, e.g., by domain decomposition.
B.2.2 Variational flow matching
To obtain further insights in how flow matching works and lay the ground for further concepts, we describe variational flow matching [44] in the following. In the end, we demonstrate that flow matching can be seen as a generalisation of the deterministic model that refines the predictions instead of making a one-step projection.
@@ B.2.2 Variational flow matching
$\mathbf{v}_{\boldsymbol{\theta}}(\mathbf{z}_{\tau},\tau)=\mathbb{E}_{\mathbf{z}_{1}\sim q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})}\left[\mathbf{u}(\mathbf{z}_{\tau}\mid\mathbf{z}_{1})\right].$ (25)
The conditional velocity $\mathbf{u}(\mathbf{z}_{\tau}\mid\mathbf{z}_{1})$ is obtained from the linear interpolation, Equation (22),
$\mathbf{u}(\mathbf{z}_{\tau}\mid\mathbf{z}_{1})=\frac{\mathbf{z}_{1}-\mathbf{z}_{\tau}}{1-\tau}.$ (26)
Since this conditional velocity is analytically defined, the approximated velocity, Equation (25), matches the true velocity, if the variational distribution matches the true distribution $p(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})$, which suggests a minimisation of the Kullback-Leibler divergence between these two distributions as loss function,
@@
$=-\mathbb{E}_{t,\mathbf{z}_{0},\mathcal{D}}\left[\log\big{(}q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})\big{)}\right]+C,$ (27)
where $C$ represents terms that do not depend on $\boldsymbol{\theta}$. Thanks to a mean-field approximation [44], the log-likelihood can further be factorised over the $n_{\mathrm{vars}}$ variables and the $n_{\mathrm{grid}}$ grid points to
$\mathcal{L}_{\textrm{VFM}}(\boldsymbol{\theta})=-\mathbb{E}_{t,\mathbf{z}_{0},\mathcal{D}}\left[\sum^{n_{\mathrm{vars}}}_{k=1}\sum^{n_{\mathrm{grid}}}_{l=1}\log\big{(}q_{\boldsymbol{\theta}}(z_{1,k,l}\mid\mathbf{z}_{\tau})\big{)}\right]+C$ (28)
@@
Variational flow matching is a generalisation of the previously stated flow matching. The linearity of the conditional velocity allows us to reduce the variational distribution to a mean-field approximation, where the distribution over all variables and grid points is factorised. By combining mean-field variational flow matching with a Gaussian assumption, we recover conventional flow matching up to a constant [44],
$q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})=\mathcal{N}\left(\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}},(1-\tau)^{2}\mathbf{I}\right)$ (29)
@@
$\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}}=\mathbf{z}_{\tau}+(1-\tau)\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h}).$ (30)
@@
$\propto\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\lVert\mathbf{v}-\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})\rVert^{2}_{2}\right],$ (31)

### v2

@@ Acknowledgments
Scaling analysis
In addition to the resolved dynamics, sea ice is subject to sub-grid dynamics that act at scales below the resolved $12\,\text{km}$ resolution with a brittle behaviour. For a qualitative analysis, we imitate how features are tracked and deformations rates are calculated with the Environment and Climate Change Canada automated sea ice tracking system (ECCC-ASITS) based on Sentinel-1 synthetic aperture radar (SAR) imagery (?) (data available at https://zenodo.org/records/13936609, last access: 5th November 2025). We use the individual feature pairs and their triangulation found by ECCC-ASITS from 2018-02-16 03:00 to 2018-02-25 15:00. Based on their temporal and spatial difference, we estimate the features’ drift by the relation
$\bar{\mathbf{u}}_{t}=\frac{\mathbf{x}_{t}-\mathbf{x}_{s}}{t-s},$ (S11)
where $\mathbf{x}_{t}$ is the position at time $t$ and $\mathbf{x}_{s}$ the position at time $s$ with $t>s$. We use the original per satellite triangulation of the features and retain all triangles where the original signal-to-noise ratio is above $0.5$. The drift gradients, needed for the deformation rates, are estimated based on the triangles with the contour integral (?), e.g., $u_{x}=\frac{1}{A}\oint u\,dy$ for the $u$-drift gradient in $x$-direction with $A$ as triangle area. The shear deformation, $\dot{\epsilon}_{\mathrm{shear}}$, divergent deformation, $\dot{\epsilon}_{\mathrm{div}}$, and total deformation rates $\dot{\epsilon}_{\mathrm{tot}}$ are estimated by
$\dot{\epsilon}_{\mathrm{shear}}$ $=\sqrt{(u_{x}-v_{y})^{2}+(u_{y}+v_{x})^{2}},$ (S12)
$\dot{\epsilon}_{\mathrm{div}}$ $=u_{x}+v_{y},$ (S13)
$\text{and }\dot{\epsilon}_{\mathrm{tot}}$ $=\sqrt{\dot{\epsilon}^{2}_{\mathrm{shear}}+\dot{\epsilon}^{2}_{\mathrm{div}}}.$ (S14)
Every 12 hours, we initialise particles at positions of the found features in a $\pm$ 6-hour window. Based on the simulated sea-ice drift in the GenSIM and neXtSIM simulations, we advect all particles forward to approximately match the same period ($\pm$ 6 hours) as used for the original features. As advection scheme, we use a simple forward Euler integration with an hourly time step, a bilinear drift interpolation in space, and linear interpolation in time. The averaged drift and deformation rates for GenSIM and neXtSIM are then estimated based on these advected particles.
GenSIM demonstrates its strong skill to reproduce key brittle characteristics of observed Arctic sea-ice deformation from Sentinel-1 SAR images, while performing comparably to neXtSIM (Fig. S5a–c). GenSIM successfully captures the spatial organisation and intensity of the deformation features from SAR and neXtSIM, including anisotropic features and high-strain bands, even though neXtSIM and GenSIM have more diffuse structures due to their spatial and temporal averaging (the predicted drift is six-hour averages).
In neXtSIM, the sub-grid dynamics are parametrised by the brittle Bingham-Maxwell rheology (?), thanks to which the model can reproduce multifractal resolved short-term dynamics as observed by satellites. To quantitatively compare the scaling of GenSIM’s and neXtSIM’s short-term dynamics in space and time, we initialise an ensemble of $65\times 65$ particles at the cell centres of the curvilinear mesh in the central Arctic, far from the coast, at 2017-12-01 03:00. As the curvilinear mesh is structured, the particles naturally form quadrangles. These particles are then continuously advected with the predicted drift from GenSIM and neXtSIM based on the same advection scheme as described before. We track the positions of these advected particles from 2017-12-01 03:00 to 2018-03-31 15:00. The averaged drift is then calculated based on the difference between two corresponding positions in time with Eq. S11.
To estimate the temporal and spatial scaling, we coarse-grain by calculating averaged drifts at various time periods and by averaging drift gradients, respectively. The drift gradients from the original quadrangles are merged into larger quadrangles with an area-weighted averaging (?). With these larger quadrangles and their averaged gradients, we estimate the deformation rates at nominal temporal scales $T\in[12\,\text{h},1\,\text{d},2\,\text{d},4\,\text{d},8\,\text{d},16\,\text{d},32\,\text{d}]$} and at nominal spatial scales $L\in[12.25\,\text{km},24.5\,\text{km},49\,\text{km},98\,\text{km},196\,\text{km},392\,\text{km}]$}. For each spatial scales, we require that $2^{-1/2}L\leq l\leq 2^{1/2}L$ holds for the area of the merged quadrangles $l$.
At each scale, we track the first three distributional moments of the deformation rates. The scaling analysis then assumes power laws in space and time such that the moments $q$ of the averaged rates at scale $L$ and $T$ follow $\big\langle\dot{\epsilon}^{q}(L,T)\big\rangle\sim L^{-\beta(q)}$ and $\big\langle\dot{\epsilon}^{q}(L,T)\big\rangle\sim T^{-\alpha(q)}$, respectively. The scaling coefficients $\beta(q)$ and $\alpha(q)$ in space and time, respectively, describe how the moments change by coarse-graining. The dependence of the scaling coefficients on the moment gives an estimate how much the localisation of fracturing events are dependent on the magnitude of these events, a positive curvature indicates thereby a multifractal and intermittent behaviour (?, ?).
Here, GenSIM maintains the behaviour of the sea-ice dynamics as modelled by neXtSIM: the heavy-tailed distribution of total deformation rates from GenSIM mirrors neXtSIM’s at the 12 km and 12 hour scale (Fig. S5d). The spatial scaling of GenSIM’s dynamics closely aligns with neXtSIM’s across orders of magnitude (Fig. S5e), even though GenSIM has a small positive bias due to an overestimation of intermediate deformation rates. Moreover, GenSIM yields a similar intermittency to neXtSIM across a range of temporal averaging windows (Fig. S5f). Despite generally matching neXtSIM, GenSIM slightly underestimates the scaling coefficients at higher moments, as indicated by the slopes of the fitted power laws for $q=3$. Consequently, GenSIM shows a multifractal spatial and temporal scaling to a slightly lesser extent than neXtSIM. Nevertheless, the scaling invariance of GenSIM is remarkable, considering that it is estimated based on diagnosed deformation rates, which require the right correlations in space and time between the predicted drift components. This underlines once again GenSIM’s physical consistency, which allows us to perform such advanced analyses.
Sea-ice extent
The sea-ice extent is evaluated similarly to the sea-ice volume in the main manuscript. All simulated and retrieved sea-ice concentration products are bilinearly interpolated to the EASE-25-km mesh from OSISAF and averaged to a monthly timescale. The monthly averaged sea-ice concentration is thresholded by the usual threshold of $\geq 0.15$.
For April and September, all products have a similar evolution of the sea-ice extent (Fig. S6a) with small positive biases in September for PIOMAS before 2007 and for GenSIM after 2007. This positive bias in GenSIM is also evident in the seasonal cycle of the extent for the period 2015–2018 (Fig. S6b). Nonetheless, GenSIM’s 20-year trend for individual months is remarkably similar to OSISAF’s. GenSIM successfully shows the slowing in the decrease of the sea-ice extent (?) (Fig. S6c–e). This becomes especially striking when compared to the other products for the 20-year trend in September. Therefore, GenSIM simulates the Arctic-wide sea-ice extent as observed, suggesting its reliability for longer climate-like projections.
Sea-ice thickness and area
As the sea-ice volume is the product of the sea-ice thickness with the sea-ice concentration, we evaluate these two components independently with the averaged sea-ice thickness and the sea-ice area. Like previously, all fields are bilinearly interpolated to the common EASE-25-km mesh and averaged to monthly values. The thickness of a grid cell is only used for the averaged thickness, if its averaged concentration is above $0.15$.
GenSIM follows the trend in the sea-ice thickness towards thinner ice remarkably well and matches neXtSIM with a small positive bias (Fig. S7a). Before 2007, both GenSIM and neXtSIM have the tendency towards thinner ice than inferred from PIOMAS. Yet, after 2007, GenSIM’s average thickness captures the annual variations in PIOMAS and the IceSAT2 satellite product. This agreement becomes especially evident in April. GenSIM and neXtSIM have both a sea-ice thickness minimum in September 2017 and 2018, which also drives their minimum for the sea-ice volume. While the small bias of GenSIM compared to neXtSIM manifests in the annual cycle, it fits the cycle of PIOMAS from August to April better (Fig. S7b). Compared to all other products, even to IceSAT2, CS2/SMOS exhibits a neagtive bias in the thickness that increases with increasing freezing period. Similar to the sea-ice extent, the averaged area of GenSIM is in the range of all other products, include the OSISAF and NSIDC climate data records (Fig. S7c&d). Hence, these results suggest that existing differences and biases are mostly due to differences in the sea-ice thickness while the sea-ice extent and area are rather well constrained.
Diagnosed sea-ice age
The sea-ice age is diagnosed as incremental tracer variable, which is advected with the predicted drift and retained over ice-covered mesh cells. To advect in Eulerian space, we use a backward semi-Lagrangian integration scheme (?, ?): all unmasked cell centres are advected $12\,\text{h}$ backward in time with an hourly time step. The predicted drift is mapped to advected centres with a nearest neighbour interpolation in space and a linear interpolation in time. After advecting the centres, each unmasked cell with its centre, $(x_{t},y_{t})$, has a corresponding displaced centre $12$ hours earlier, $(x^{\prime}_{t-12\,\text{h}},y^{\prime}_{t-12\,\text{h}})$. To obtain the advected sea-ice age, we bilinearly interpolate the gridded sea-ice age, $\mathrm{age}_{t-12\,\text{h}}(x^{\prime}_{t-12\,\text{h}},y^{\prime}_{t-12\,\text{h}})$. The new sea-ice age reads then
$\mathrm{age}_{t}(x_{t},y_{t})=$ $\begin{cases}\mathrm{age}_{t-12\,\text{h}}(x^{\prime}_{t-12\,\text{h}},y^{\prime}_{t-12\,\text{h}})+12\,\text{h},&\text{if }\mathrm{conc}_{t}(x_{t},y_{t})>0.15\\ 0,&\text{otherwise}.\end{cases}$ (S15)
Effectively, the sea-ice age is reset if the mesh cell is no longer covered by ice (concentration <0.15<0.15) or if the advected age comes from land. Caused by the bilinear interpolation within the advection scheme, the predicted sea-ice age is exposed to some diffusion.
To assess the impact of the warming climate on the predicted sea-ice age, we advect the sea-ice age from 2000-01-01 03:00 to 2003-12-31 15:00 and from 2015-01-01 03:00 to 2018-12-31 15:00 for GenSIM and neXtSIM. Hence, the upper bound for the sea-ice age at the end of these windows is four years. These estimates are compared to the satellite sea-ice age product from the National Snow and Ice Data Center (NSIDC) (?) on the EASE-12.5-km mesh.
Remarkably, GenSIM, neXtSIM and the satellite product agree quite well in the large-scale distribution of the sea-ice age for 2003 (Fig. S8a–c), as well as 2018 (Fig. S8d–f). Also, all products show a transition from large fractions of multi-year ice in 2003 to larger fractions in single-year ice or younger in 2018. To mirror the observed sea-ice age, the models needs to represent the correct evolution and correlations in the sea-ice drift and sea-ice concentration as similarly observed. Hence, GenSIM can represent them like neXtSIM does, suggesting that GenSIM performs simulations that have similar quality as those from neXtSIM.
Generalisation to other resolutions
GenSIM is designed to work in resolutions between 12 km and 100 km by randomly averaging the data during training. To test the capabilities of GenSIM at other resolutions than the original 12 km, we perform an experiment for ten years at 96 km. Initialised with a $8\times 8$ coarse-grained predictions from the high-resolution GenSIM run on 2015-01-01 03:00, the model is run until 2024-12-31 15:00 with coarse-grained ERA5 forcings. This low-resolution simulation is compared to the high-resolution simulation and the PIOMAS reanalysis.
For all three products, the sea-ice volume is estimated by bilinearly interpolating onto a common EASE-100-km mesh. As for the original volume estimation, a common mask is used and all data is monthly averaged. Please note that the sea-ice volume estimates are below the ones from Fig. 3, as a coarser masking has been used.
The low-resolution simulation follows remarkably closely the previously established performance of GenSIM at high resolutions (Fig. S9a). The bias of GenSIM at high resolutions compared to PIOMAS is slightly reduced for the low-resolution simulation, while the seasonal cycle is matched almost perfectly. The representation of thermodynamics in GenSIM apparently works even at coarse resolutions, and we would expect similar results at other resolutions within the trained scales
We analyse the dynamics with the same prediction time as for Fig 2 by estimating the divergent deformation rate on 2024-02-13 03:00. In addition to estimating the deformation rate at the high resolution (Fig. S9b) and low resolution (Fig. S9d), we also coarse-grain the drift from the high resolution to the low resolution, with which we then estimate the deformation rates (Fig. S9c).
By coarse-graining the drift from the high resolution, we attenuate the deformation rates and average out small-scale variations such that only large-scale patterns remain. However, when run at low resolutions, GenSIM can suddenly resolve features from the high-resolution simulation which were strongly averaged out during the coarse graining, e.g., the divergent lead in the crossing the central Arctic from the south-east to the north-west. Consequently, at the low resolution, GenSIM exhibits dynamics as we would similar expect them to behave if the sub-grid scales are properly parametrised.
These results suggest that GenSIM learns to generalise its dynamics and thermodynamics at different resolutions. Hence, the strategy of training the model with randomly coarse-grained data was successful and is apparently a way to artificially increase the data amount. Furthermore, this renders GenSIM as a highly flexible sea-ice model, indicating its potential to replace sea-ice component models in coupled Earth system models.
Mathematical framework for surrogate modelling
@@
Our goal is to develop a powerful surrogate model that forecasts sea-ice states with a 12-hour lead time, $\mathbf{x}_{t+12\,\text{h}}$, given some initial states $\mathbf{x}_{t}$ and external forcings $\mathbf{F}_{t:t+12\,\text{h}}$ provided at initial time $t$ and 12 hours later, $t+12\,\text{h}$. As a reminder, the predictive variables are the sea-ice thickness, the concentration, the damage, the two drift components, and the snow-on-ice thickness. Our model predicts the tendency within the 12 hours,
$\mathbf{x}_{t+12\,\text{h}}\approx\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}}=\mathbf{x}_{t}+\mathcal{M}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\mathbf{F}_{t:t+12\,\text{h}}),$ (S16)
@@
with $\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}}$ as the model prediction. The tendency model $\mathcal{M}_{\boldsymbol{\theta}}(\cdot)$ is based on a deep neural network with its learnable weights and biases $\boldsymbol{\theta}$. Here, we train two different types of models: a deterministic model (?) and a generative model based on flow matching (?, ?, ?).
@@
To train the models, we sample data triplets from a training dataset $(\mathbf{x}_{t},\mathbf{x}_{t+12\,\text{h}},\mathbf{F}_{t:t+12\,\text{h}})\sim\mathcal{D}$ and minimise the loss functions using variants of stochastic gradient descent. Once trained, we employ the models as outlined in Equation S16 to propagate the initial states forward and to generate a 12-hour lead time prediction. For predictions extending beyond 12 hours, we utilise the model autoregressively, wherein the preceding prediction serves as the input for the subsequent time step, with the forcings being updated accordingly.
@@
$\mathbf{z}=\frac{\mathbf{x}_{t+12\,\text{h}}-\mathbf{x}_{t}}{\boldsymbol{\sigma}_{\textrm{clim}}}.$ (S17)
@@
$\mathbf{x}_{t+12\,\text{h}}\sim\mathcal{N}\left(\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}},\boldsymbol{\sigma}_{\textrm{clim}}^{2}\mathbf{I}\right),$ (S18)
@@
$\mathcal{L}_{\textrm{det}}(\boldsymbol{\theta})=\mathbb{E}_{\mathcal{D}}\left[\mathbf{w}\lVert\mathbf{x}_{t+12\,\text{h}}-\widehat{\mathbf{x}}_{\boldsymbol{\theta},t+12\,\text{h}}\rVert^{2}_{2}\right],$ (S19)
with $\mathbf{w}=1/{\boldsymbol{\sigma}_{\textrm{clim}}^{2}}$ as weighting vector. This loss function is commonly used for other surrogate models (?, ?, ?) and serves as a baseline for our generative model.
@@
$\mathcal{L}_{\textrm{det}}(\boldsymbol{\theta})=\mathbb{E}_{\mathcal{D}}\left[\lVert\mathbf{z}-\widehat{\mathbf{z}}_{\boldsymbol{\theta}}\rVert^{2}_{2}\right].$ (S20)
@@
$\mathbf{z}_{0}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}).$ (S21)
@@
$\frac{\mathrm{d}\mathbf{z}_{\tau}}{\mathrm{d}\tau}=\mathbf{u}(\mathbf{z}_{\tau}),$ (S22)
@@
$\mathbf{z}_{\tau}=\tau\mathbf{z}_{1}+(1-\tau)\mathbf{z}_{0},$ (S23)
@@
$\mathbf{u}=\mathbf{z}_{1}-\mathbf{z}_{0}.$ (S24)
Since $\mathbf{z}_{1}$ is targetted and the analytical velocity therefore unknown during prediction, we train the neural network with its parameters $\boldsymbol{\theta}$ to approximate the true velocity, $\mathbf{u}\approx\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})$. By conditioning the approximation on the initial states and forcings, the neural network learns to sample from the conditional distribution (?).
@@
$\mathcal{L}_{\textrm{FM}}(\boldsymbol{\theta})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\lVert\mathbf{u}-\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z}_{\tau},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})\rVert^{2}_{2}\right].$ (S25)
Once trained, we can employ the neural network as velocity approximation for the ODE, Equation S22. By initialising with Equation S21 and integrating the ODE in pseudo time from $\tau=0$ to $\tau=1$, we produce a sample from the target distribution, conditioned on the initial states and forcings. By drawing different initialisations, an ensemble of predictions can be produced from the same conditional information. In practise, we perform the integration with a second-order Heun scheme (?) in 20 steps, where the discrete steps are scheduled with a shifted sigmoid function, as shown in Fig. S10b.
Notably, if projected back into physical space, the generative flow is initialised by a sample from the prior distribution around the initial conditions $\mathcal{N}(\mathbf{x}_{t},\boldsymbol{\sigma}^{2}_{\text{clim}}\mathbf{I})$ and maps to a sample from the predictive distribution for $\mathbf{x}_{t+12\,\text{h}}$. Although a deterministic forecast, trained with Equation S19, could also be used to define the prior (?, ?), previous work (?) has demonstrated negligible performance difference, and we therefore employ a persistence forecast as prior for simplicity. Since the integration of the ODE maps from the initial conditions to the prediction, it works similarly to classical geophysical models (?), integrating in pseudo time instead of real time. Hence, we can make use of established numerical tools to render generative flow models more efficient and to scale them up, e.g., by domain decomposition.
To obtain further insights in how flow matching works and lay the ground for further concepts, we describe variational flow matching (?) in the following. In the end, we demonstrate that flow matching can be seen as a generalisation of the deterministic model that refines the predictions instead of making a one-step projection.
@@
$\mathbf{v}_{\boldsymbol{\theta}}(\mathbf{z}_{\tau},\tau)=\mathbb{E}_{\mathbf{z}_{1}\sim q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})}\left[\mathbf{u}(\mathbf{z}_{\tau}\mid\mathbf{z}_{1})\right].$ (S26)
The conditional velocity $\mathbf{u}(\mathbf{z}_{\tau}\mid\mathbf{z}_{1})$ is obtained from the linear interpolation, Equation S23,
$\mathbf{u}(\mathbf{z}_{\tau}\mid\mathbf{z}_{1})=\frac{\mathbf{z}_{1}-\mathbf{z}_{\tau}}{1-\tau}.$ (S27)
Since this conditional velocity is analytically defined, the approximated velocity, Equation S26, matches the true velocity, if the variational distribution matches the true distribution $p(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})$, which suggests a minimisation of the Kullback-Leibler divergence between these two distributions as loss function,
@@
$=-\mathbb{E}_{t,\mathbf{z}_{0},\mathcal{D}}\left[\log\big(q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})\big)\right]+C,$ (S28)
where $C$ represents terms that do not depend on $\boldsymbol{\theta}$. Thanks to a mean-field approximation (?), the log-likelihood can further be factorised over the $n_{\mathrm{vars}}$ variables and the $n_{\mathrm{grid}}$ grid points to
$\mathcal{L}_{\textrm{VFM}}(\boldsymbol{\theta})=-\mathbb{E}_{t,\mathbf{z}_{0},\mathcal{D}}\left[\sum^{n_{\mathrm{vars}}}_{k=1}\sum^{n_{\mathrm{grid}}}_{l=1}\log\big(q_{\boldsymbol{\theta}}(z_{1,k,l}\mid\mathbf{z}_{\tau})\big)\right]+C$ (S29)
@@
Variational flow matching is a generalisation of the previously stated flow matching. The linearity of the conditional velocity allows us to reduce the variational distribution to a mean-field approximation, where the distribution over all variables and grid points is factorised. By combining mean-field variational flow matching with a Gaussian assumption, we recover conventional flow matching up to a constant (?),
$q_{\boldsymbol{\theta}}(\mathbf{z}_{1}\mid\mathbf{z}_{\tau})=\mathcal{N}\left(\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}},(1-\tau)^{2}\mathbf{I}\right)$ (S30)
@@
$\widehat{\boldsymbol{\mu}}_{\boldsymbol{\theta}}=\mathbf{z}_{\tau}+(1-\tau)\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h}).$ (S31)
@@
$\propto\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\lVert\mathbf{v}-\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z},\tau,\mathbf{x}_{t},\mathbf{F}_{t:t+12h})\rVert^{2}_{2}\right],$ (S32)